# Two dimensional dimers beyond planarity

### v2

@@ 7.1 Multiple dimer model
[…]
obtaining that the RHS of (8.1) equals
@@
$\sum\limits_{m^{1}\in\mathbb{N}_{0}^{E}}\sum\limits_{m^{2}\in\mathbb{N}_{0}^{E}}\sum\limits_{m^{g,1}\in(r\mathbb{N}_{0})^{E_{g}}}\sum\limits_{m^{g,2}\in(r\mathbb{N}_{0})^{E_{g}}}\prod_{e\in E}\frac{\beta^{m^{1}_{e}}}{m^{1}_{e}!}\prod_{e\in E}\frac{\beta^{m^{2}_{e}}}{m^{2}_{e}!}\prod_{e\in E_{g}}\frac{(\beta h)^{\frac{m^{g,1}_{e}}{r}}}{\frac{m^{g,1}_{e}}{r}!}\prod_{e\in E_{g}}\frac{(\beta h)^{\frac{m^{g,2}_{e}}{r}}}{\frac{m^{g,2}_{e}}{r}!}\\ \times\int_{S^{V}}\boldsymbol{ds}\,\boldsymbol{g({s})}\prod_{z\in A}g_{u,z}(s_{z})\prod_{x\in V^{e}}e^{i(u^{1}_{x}s_{x}^{1}-u^{2}_{x}s_{x}^{2}+\,\tilde{n}_{x}(m^{1})s_{x}^{1}\,-\tilde{n}_{x}(m^{2})s_{x}^{2}+m_{\{x,g\}}^{g,1}s_{x}^{1}-m_{\{x,g\}}^{g,2}s_{x}^{1})}\\ \prod_{x\in V^{o}}e^{-i(u^{1}_{x}s_{x}^{1}-u^{2}_{x}s_{x}^{2}+\,\tilde{n}_{x}(m^{1})s_{x}^{1}\,-\tilde{n}_{x}(m^{2})s_{x}^{2}+m_{\{x,g\}}^{g,1}s_{x}^{1}-m_{\{x,g\}}^{g,2}s_{x}^{1})},$ (8.2)
where we set $m^{g,2}_{\{x,g\}}=m^{g}_{\{x,g\}}-m^{g,1}_{\{x,g\}}$, and interpret $m^{g,i}_{\{x,g\}}$ as the number of dimers of colour $i$ on $\{x,g\}$. Given $n_{1},n_{2}\in\mathbb{Z}$ we use for each vertex in $V\setminus A$ the identity
$(\frac{1}{2\pi})^{2}\int_{S}e^{i(n_{1}s^{1}-n_{2}s^{2})}\sum\limits_{n=0}^{\infty}U(n)e^{-i\,n\,(s^{1}-s^{2})}ds=U(n_{1})\mathbbm{1}_{\{n_{1}=n_{2}\}},$
and for each vertex in $A$ the identity
[…]
obtaining that, for any fixed $m^{1},m^{2},m^{1,g}$, $m^{2,g}$ in the sum (8.2), the corresponding integral equals
$\prod_{x\in V\setminus A}\mathbbm{1}_{\{\tilde{n}_{x}(m^{1})+m_{\{x,g\}}^{g,1}+u_{x}^{1}=\tilde{n}_{x}(m^{2})+u_{x}^{2}+m_{\{x,g\}}^{g,2}\}}U(\tilde{n}_{x}(m^{1})+u_{x}^{1}+m_{\{x,g\}}^{g,1}-m_{\{x,g\}}^{g,2})\\ \times\prod_{x\in A}\mathbbm{1}_{\{\tilde{n}_{x}(m^{1})+m_{\{x,g\}}^{g,1}+u_{x}^{1}-m_{\{x,g\}}^{g,2}=\tilde{n}_{x}(m^{2})+u_{x}^{2}=u\}}U(u).$ (8.3)
For $x\in V$, we further write $u_{x}^{i}=(u_{+}^{i})_{x}+(u_{-}^{i})_{x}$ (recalling the definitions in the statement of the proposition) and interpret $(u_{+}^{1})_{x}-(u_{-}^{2})_{x}$ as the number of $1$-dimers on $\{x,s\}$ and $-(u_{-}^{1})_{x}+(u_{+}^{2})_{x}$ as the number of $2$-dimers on $\{x,s\}$ (so that these numbers are always non negative). Thus, $n_{x}^{1}=n_{x}^{1}(m^{1},m^{1,g},u^{1},u^{2})=\tilde{n}_{x}(m^{1})+m_{\{x,g\}}^{1,g}+(u^{1}_{+})_{x}-(u^{2}_{-})_{x}$ corresponds to the number of $1$-dimers touching $x\in V$ and $n_{x}^{2}=n_{x}^{2}(m^{2},m^{2,g},u^{1},u^{2})=\tilde{n}^{2}_{x}(m^{2})+m_{x,g}^{2,g}-(u^{1}_{-})_{x}+(u^{2}_{+})_{x}$ corresponds to the number of $2$-dimers touching $x\in V$. We then obtain that, for any pair $m=(m^{1},m^{2})\in\mathbb{N}_{0}^{E_{en}}\times\mathbb{N}_{0}^{E_{en}}$, the previous expression can be reformulated as follows
$\prod_{x\in V\setminus A}\mathbbm{1}_{\{n_{x}^{1}(m)=n_{x}^{2}(m)\}}U\big{(}n^{1}_{x}(m)-\,m_{\{x,g\}}^{g,2}+(u_{-}^{1})_{x}+(u_{-}^{2})_{x}\big{)}\prod_{x\in A}\mathbbm{1}_{\{n_{x}^{1}(m)=n_{x}^{2}(m)=u+m_{\{x,g\}}^{g,2}-(u_{-}^{1})_{x}-(u_{-}^{2})_{x}\}}U(u).$ (8.4)
Recalling the definition of $\Sigma$ and the notation (3.1) we then obtain that the LHS of (8.1) equals
@@
$\sum\limits_{\begin{subarray}{c}m=(m^{1},m^{2})\in\Sigma:\\ \partial m=(u^{1}_{+}-u^{2}_{-},\,u^{2}_{+}-u^{1}_{-})\\ n_{z}(m)=m_{\{z,g\}}^{2}-(u_{-}^{1})_{z}-(u_{-}^{2})_{z}+u\,\,\,\forall z\in A\end{subarray}}\tilde{\mu}_{\beta,h,r}(m)\,\,\prod_{x\in V}U\big{(}n_{x}(m)-\,m_{\{x,g\}}^{2}+(u_{-}^{1})_{x}+(u_{-}^{2})_{x}\big{)}.$
By further observing that for each $m\in\Sigma$ we have precisely $n_{x}(m)!$ possible pairings at each vertex $x\in V$, we deduce (3.9) and (3.10). We also see that if we replace $g_{u,z}$ with $g_{z}$ in the product in (8.1) we obtain that the sum over $m^{g}\in\mathbb{N}_{0}^{E_{g}}$ in the next expression replaced by a sum over $m^{g}\in\mathbb{N}_{0}^{E_{g}}$ such that $m_{\{x,g\}}=0$ for each $x\in A$ (i.e, no dimer on $\{x,g\}$ for $x\in A$ is present), by further assuming that $u_{x}^{i}=0$ for each $x\in V$ we then obtain that (8.4) is replaced by
$\prod_{x\in V\setminus A}\mathbbm{1}_{\{n_{x}^{1}(m)=n_{x}^{2}(m)\}}U\big{(}n^{1}_{x}(m)\big{)}\prod_{x\in A}\mathbbm{1}_{\{n_{x}^{1}(m)=n_{x}^{2}(m)=0\}}U(0),$ (8.5)
this leads to (3.11) following the same steps as for the previous display. ∎

### v3

@@ 8.1 Multiple-type dimer model
We define for each $\boldsymbol{s}\in\Omega_{s}$ the function
$\gamma_{z}(\boldsymbol{s}):=\frac{1}{(2\pi)^{2N}}\sum\limits_{n_{1},\ldots,n_{N}\geq 0}^{\infty}U(n_{1},\ldots,n_{N})\prod_{i=1}^{N}\big(\overline{S_{z}^{2i-1}(\boldsymbol{s})\,\,S_{z}^{2i}(\boldsymbol{s})}\big)^{n_{i}},$ (8.2)
$U:\mathbb{N}_{0}^{N}\mapsto\mathbb{R}_{0}^{+}$ generalizes the weight function, and, as before, $\boldsymbol{\gamma}(\boldsymbol{s}):=\prod_{z\in V}\gamma_{z}(\boldsymbol{s}).$ For each $\boldsymbol{s}\in\Omega_{s}$ the Hamiltonian function, $H:\Omega_{s}\mapsto\mathbb{C}$, is defined as
$H(\boldsymbol{s})=\sum\limits_{i=1}^{2N}\sum\limits_{\{x,y\}\in E}S_{x}^{i}(\boldsymbol{s})S_{y}^{i}(\boldsymbol{s})+h\sum\limits_{z\in V}\,2\,\cos(rs^{1}_{z}),$
By adapting the proof of Proposition 3.11 to this setting we deduce that such a spin system corresponds to a multi-occupancy double-dimer model which is similar to the one which has been introduced in Section 3. The only difference is that dimers have $2N$ possible ‘colours’, which we denote by $1,\ldots,2N$, and, similar to the model which was introduced in Section 3, the number of dimers of odd colour $i$ equals the number of dimers of colour $i+1$ at each vertex and for each $i$. In the special case
$U(n_{1},n_{2},\ldots n_{N})=\begin{cases}1&\mbox{ if $n_{1}+\ldots+n_{N}\in\{0,1\}$,}\\ 0&\mbox{ otherwise,}\end{cases}$
for example, one obtains a generalization of the monomer double-dimer model in which each vertex is either isolated or it belongs to a sequence of dimers whose colours are $2i$, 2i12i-1, $2i$, 2i12i-1, $\ldots,2i-1$ or 2i12i-1, $2i$, 2i12i-1, $2i$, $\ldots,2i$ for some integer $i\in\{1,\ldots,N\}$. Equivalently, this model can be defined by assigning to each configuration in $\Omega$, the set of monomer double-dimer configurations introduced in Section 1.3.2, the weight
$\forall\omega\in\Omega\qquad\mathbb{P}_{G,\rho}(\omega):=\frac{\rho^{|M|}N^{\mathcal{L}(\omega)}}{\mathbb{Z}_{G,N,\rho}},$ (8.3)
in place of (1.4), where $\mathcal{L}(\omega)$ is the total number of loops of at least two steps in $\omega$, and $\mathbb{Z}_{G,N,\rho}$ is a normalizing constant, giving (1.4) in the special case $N=1$.
8.2 Proof of Propositions 3.11 and 3.12
The proofs are based on the next simple identity, namely for each vertex $x\in V$, and integers $k_{1},n_{1},k_{2},n_{2}\in\mathbb{Z}$ we have
$\frac{1}{(2\pi)^{2\,|V|}}\int_{\Omega_{s}}\boldsymbol{ds}\ (\overline{S^{1}_{x}(\boldsymbol{s})})^{k_{1}}(S^{1}_{x}(\boldsymbol{s}))^{n_{1}}\ (\overline{S^{2}_{x}(\boldsymbol{s})})^{k_{2}}(S^{2}_{x}(\boldsymbol{s}))^{n_{2}}=\delta_{k_{1},n_{1}}\delta_{k_{2},n_{2}},$ (8.4)
where $\delta_{k,n}$ equals one if $k=n$ and zero otherwise. This follows from the definition of the spin variable (2.1). By using the definition of $\boldsymbol{\gamma}(\boldsymbol{s})$ (see equation (2.3)), we obtain from the previous identity that for any pair of vectors $u^{i}=(u^{i}_{x})_{x\in V}\in\mathbb{Z}^{V}$, with $i\in\{1,2\}$, the following key identity holds
$\begin{split}\int_{\Omega_{s}}\boldsymbol{ds}\,\boldsymbol{\gamma}(\boldsymbol{s})\prod_{x\in V}\prod_{i=1}^{2}(S_{x}^{i}(\boldsymbol{s}))^{{{u}^{i}_{x}}}&=\frac{1}{(2\pi)^{2\,|V|}}\sum\limits_{\boldsymbol{n}\in{\mathbb{N}_{0}}^{V}}\int_{\Omega_{s}}\boldsymbol{ds}\prod_{x\in V}U(n_{x}){\big(S_{x}^{1}(\boldsymbol{s})\big)}^{{{u}^{1}_{x}}}{\big(\overline{S_{x}^{1}(\boldsymbol{s})}\big)}^{{{n}_{x}}}{\big(S_{x}^{2}(\boldsymbol{s})\big)}^{{{u}^{2}_{x}}}{\big(\overline{S_{x}^{2}{(\boldsymbol{s})}}\big)}^{{{n}_{x}}}\\ &=\prod_{x\in V}U({{u}_{x}^{1}})\,\,\delta_{{{u}^{1}_{x}},{{u}^{2}_{x}}}\end{split}$ (8.5)
where we extended the definition of $U$ to the negative integers by setting $U(n)=0$ for each $n<0$ (this implies that the previous quantity equals zero as long as there exists some $x$ and $i\in\{1,2\}$ with $u_{x}^{i}<0$. We are now ready to prove the propositions using the previous identity. Let $f:\Omega_{s}\mapsto\mathbb{C}$ be an arbitrary function. Recall the definition of $G=(V,E)$ , of the extended graph $G_{en}=(V_{en},E_{en})$, of the set of edges incident to the source vertex, $E_{s}$, and of the set of edges incident to the ghost vertex, $E_{g}$. We observe that the factor $e^{\beta H(\boldsymbol{s})}$ in (2.5) can be seen as a product of exponential functions. Each such exponential function is associated to an edge of the extended graph and to a colour. We start by performing a Taylor expansion for each such exponential in the product obtaining
@@
$\int_{\Omega_{s}}\boldsymbol{ds}\,\boldsymbol{\gamma}(\boldsymbol{s})\,f(\boldsymbol{s})\,e^{\beta H(\boldsymbol{s})}\,=\sum\limits_{m^{1}\in\mathbb{N}_{0}^{E}}\prod_{e\in E}\frac{\beta^{m^{1}_{e}}}{m^{1}_{e}!}\sum\limits_{m^{2}\in\mathbb{N}_{0}^{E}}\prod_{e\in E}\frac{\beta^{m^{2}_{e}}}{m^{2}_{e}!}\sum\limits_{m^{g}\in\mathbb{N}_{0}^{E_{g}}}\prod_{e\in E_{g}}\frac{(\beta h)^{m^{g}_{e}}}{m^{g}_{e}!}\\ \times\int_{\Omega_{s}}\boldsymbol{ds}\,\boldsymbol{\gamma}(\boldsymbol{s})f(\boldsymbol{s})\prod_{x\in V}\Big((S_{x}^{1}(\boldsymbol{s}))^{r}+(\overline{S_{x}^{1}(\boldsymbol{s})})^{r}\Big)^{m_{\{x,g\}}^{g}}\prod_{x\in V}{(S_{x}^{1}(\boldsymbol{s}))}^{\tilde{n}_{x}(m^{1})}{(S_{x}^{2}(\boldsymbol{s}))}^{\tilde{n}_{x}(m^{2})}$ (8.6)
where we used that $2\cos(rs_{x}^{1})=(S_{x}^{1}(\boldsymbol{s}))^{r}+(\overline{S_{x}^{1}(\boldsymbol{s})})^{r}$, we defined ${\tilde{n}_{x}}(m^{i})=\sum_{y\sim x}m_{\{x,y\}}^{i}$, and $y\sim x$ means that $y$ and $x$ are neighbour in the original graph $G$ (this neighbourhood relation, then, does not consider the edges incident to the ghost or source vertex). In the previous sum we interpret $m_{e}^{i}$ as the number of dimers of colour $i$ on the edge $e$ and $m_{\{x,g\}}$ as number of dimers on the edge $\{x,g\}$ (independently from their colour). We now use at each vertex $x\in V$ the following identity
$\sum\limits_{m_{\{x,g\}}=0}^{\infty}\frac{(\beta h)^{m_{\{x,g\}}}}{m_{\{x,g\}}!}\Big((S_{x}^{1})^{r}+(\overline{S_{x}^{1}})^{r}\Big)^{m_{\{x,g\}}}$ $=\sum\limits_{m_{\{x,g\}}^{1}=0}^{\infty}\sum\limits_{m_{\{x,g\}}^{2}=0}^{\infty}\frac{(\beta h)^{m_{\{x,g\}}^{1}}}{m^{1}_{\{x,g\}}!}\frac{(\beta h)^{m_{\{x,g\}}^{2}}}{m^{2}_{\{x,g\}}!}(S_{x}^{1})^{r\,m^{1}_{\{x,g\}}}{(\overline{S_{x}^{1}})}^{r\,m_{\{x,g\}}^{2}}$
$=\sum\limits_{m_{\{x,g\}}^{1},m_{\{x,g\}}^{2}\in r\mathbb{N}_{0}}\frac{(\beta h)^{m_{\{x,g\}}^{1}/r}}{(m^{1}_{\{x,g\}}/r)!}\frac{(\beta h)^{m_{\{x,g\}}^{2}/r}}{(m^{2}_{\{x,g\}}/r)!}(S_{x}^{1})^{\,m^{1}_{\{x,g\}}}{(\overline{S_{x}^{1}})}^{m_{\{x,g\}}^{2}}.$
and interpret $m^{1}_{\{x,g\}}$ and $m^{2}_{\{x,g\}}$ as the number of dimers of type respectively $1$ and $2$ on the edge $\{x,g\}$. Using the previous identity we then obtain that the RHS of (8.6) equals
@@
$\sum\limits_{\begin{subarray}{c}m=(m^{1},m^{2})\in(\mathbb{N}_{0}^{E\cup E_{g}})^{2}:\\ m_{\{x,g\}}^{1},m_{\{x,g\}}^{2}\in r\mathbb{N}_{0}\forall x\end{subarray}}\prod_{i=1}^{2}\prod_{e\in E}\frac{\beta^{m^{i}_{e}}}{m^{i}_{e}!}\prod_{e\in E_{g}}\frac{(\beta h)^{(m^{i}_{e}/r)}}{(m^{i}_{e}/r)!}\\ \times\int_{\Omega_{s}}\boldsymbol{ds}\,\boldsymbol{\gamma({s})}f(\boldsymbol{s})\prod_{x\in V}{(S_{x}^{1}(\boldsymbol{s}))}^{\tilde{n}_{x}(m^{1})+m_{\{x,g\}}^{1}-m_{\{x,g\}}^{2}}{(S_{x}^{2}(\boldsymbol{s}))}^{\tilde{n}_{x}(m^{2})}$ (8.7)
We are now ready to conclude the proof of the two propositions. For a lighter notation we will omit the dependence of $S_{x}^{i}$ on $\boldsymbol{s}$.
Proof of Proposition 3.11.
We let $u^{1},u^{2}\in\mathbb{Z}^{V}$ be any two vectors. We set $f=\prod_{x\in V}\prod_{i=1}^{2}(S_{x}^{i})^{u_{x}^{i}}.$ Under this choice we have that the starting expression, the LHS of (8.6), equals
$\sum\limits_{\begin{subarray}{c}m=(m^{1},m^{2})\in(\mathbb{N}_{0}^{E\cup E_{g}})^{2}:\\ m_{\{x,g\}}^{1},m_{\{x,g\}}^{2}\in r\mathbb{N}_{0}\forall x\end{subarray}}\prod_{i=1}^{2}\prod_{e\in E}\frac{\beta^{m^{i}_{e}}}{m^{i}_{e}!}\prod_{e\in E_{g}}\frac{(\beta h)^{(m^{i}_{e}/r)}}{(m^{i}_{e}/r)!}\\ \times\int_{\Omega_{s}}\boldsymbol{ds}\,\boldsymbol{\gamma({s})}\prod_{x\in V}{(S_{x}^{1})}^{u_{x}^{1}+{\tilde{n}_{x}}(m^{1})+m^{1}_{\{x,g\}}-m^{2}_{\{x,g\}}}{(S_{x}^{2})}^{u_{x}^{2}+{\tilde{n}_{x}}(m^{2})}$ (8.8)
From (8.5) we deduce that, for each $m=(m^{1},m^{2})\in\mathbb{N}_{0}^{E\cup E_{g}}$ in the previous sum,
$\begin{split}\int_{\Omega_{s}}\boldsymbol{ds}\,\boldsymbol{\gamma({s})}&\prod_{x\in V}{(S_{x}^{1})}^{u_{x}^{1}+{\tilde{n}_{x}}(m^{1})+m^{1}_{\{x,g\}}-m^{2}_{\{x,g\}}}{(S_{x}^{2})}^{u_{x}^{2}+{\tilde{n}_{x}}(m^{2})}\\ =&\prod_{x\in V}U(u_{x}^{1}+{\tilde{n}_{x}}(m^{1})+m^{1}_{\{x,g\}}-m^{2}_{\{x,g\}})\,\delta_{u_{x}^{1}+{\tilde{n}_{x}}(m^{1})+m^{1}_{\{x,g\}}-m^{2}_{\{x,g\}},u_{x}^{2}+{\tilde{n}_{x}}(m^{2})}\\ \\ =&\prod_{x\in V}U(u_{x}^{1}+{\tilde{n}_{x}}(m^{1})+m^{1}_{\{x,g\}}-m^{2}_{\{x,g\}})\,\delta_{u_{x}^{1,+}-u^{2,-}_{x}+{\tilde{n}_{x}}(m^{1})+m^{1}_{\{x,g\}},u_{x}^{2,+}-u_{x}^{1,-}+{\tilde{n}_{x}}(m^{2})+m^{2}_{\{x,g\}}},\end{split}$ (8.9)
where the components $u_{x}^{i,\pm}$ have been defined in the statement of the proposition. We see that, by (8.9), only the vectors $m=(m^{1},m^{2})\in\mathbb{N}_{0}^{E\cup E_{g}}$ such that $u_{x}^{1,+}-u^{2,-}_{x}+{\tilde{n}_{x}}(m^{1})+m^{1}_{\{x,g\}}=u_{x}^{2,+}-u_{x}^{1,-}+{\tilde{n}_{x}}(m^{2})+m^{2}_{\{x,g\}}$ at each vertex $x\in V$ are allowed to have non zero weight in the sum (8.8). We can then associate to each such vector an element $m^{\prime}$ in the dimer cardinality set $\Sigma_{r}$, which is defined in Definition 3.2, such that ${m^{\prime}}^{i}_{e}=m^{i}_{e}$ for each $e\in E\cup E_{g}$ and each $i\in\{1,2\}$ and such that ${{m^{\prime}}^{i}_{\{x,s\}}}=u_{x}^{i,+}-u^{i+1,-}$ for each $x\in V$ and $i\in\{1,2\}$ (with the convention that $i+1$ is understood modulo 2, that is, $i+1=2$ if $i=1$ and $i+1=1$ if $i=2$), so that
$U(u_{x}^{1}+{\tilde{n}_{x}}(m^{1})+m^{1}_{\{x,g\}}-m^{2}_{\{x,g\}})=U(u_{x}^{1,-}+u_{x}^{2,-}+{\tilde{n}_{x}}({m^{1}}^{\prime})+{m^{1}}^{\prime}_{\{x,g\}}-{m^{2}}^{\prime}_{\{x,g\}})$
Still denoting $m^{\prime}$ by $m$, the expression (8.8) can then be expressed as a sum over $\Sigma_{r}$ and equals then the following expression
$\sum\limits_{\begin{subarray}{c}m\in\Sigma_{r}\,:\\ m_{\{x,s\}}^{i}=u_{x}^{i,+}-u_{x}^{i+1,-}\,\,\forall x\in V,\,\,\forall i\in\{1,2\}\end{subarray}}\prod_{i=1}^{2}\prod_{e\in E}\frac{\beta^{m^{i}_{e}}}{m^{i}_{e}!}\prod_{e\in E_{g}}\frac{(\beta h)^{m^{i}_{e}/r}}{(m^{i}_{e}/r)!}\prod_{x\in V}U({\tilde{n}_{x}}(m^{1})+m^{1}_{\{x,g\}}-m^{2}_{\{x,g\}}+u_{x}^{1,-}+u_{x}^{2,-}).$ (8.10)
This corresponds to the numerator in the RHS of (3.11) and concludes the proof. ∎
Proof of Proposition 3.12.
We set $f=\prod_{z\in A}\frac{U(0)\,e^{-2h\beta\cos(rs_{z}^{1})}}{(2\pi)^{2}\,\gamma_{z}(\boldsymbol{s})}$ and obtain from (8.7) that the starting expression, the LHS of (8.6), under this choice of $f$ equals
$\sum\limits_{\begin{subarray}{c}m=(m^{1},m^{2})\in\mathbb{N}_{0}^{E\cup E_{g}}:\\ m^{1}_{\{x,g\}}=m^{2}_{\{x,g\}}=0\,\,\forall x\in A\\ m_{\{x,g\}}^{1},m_{\{x,g\}}^{2}\in r\mathbb{N}_{0}\,\,\forall x\in V\end{subarray}}\prod_{i=1}^{2}\prod_{e\in E}\frac{\beta^{m^{i}_{e}}}{m^{i}_{e}!}\prod_{e\in E_{g}}\frac{(\beta h)^{m^{i}_{e}/r}}{(m^{i}_{e}/r)!}\\ \times\int_{\Omega_{s}}\boldsymbol{ds}\,\big(\frac{U(0)}{(2\pi)^{2}}\big)^{|A|}\prod_{z\in V\setminus A}\,{\gamma_{z}(\boldsymbol{s})}\,\prod_{x\in V}{(S_{x}^{1})}^{{\tilde{n}_{x}}(m^{1})+m^{1}_{\{x,g\}}-m^{2}_{\{x,g\}}}{(S_{x}^{2})}^{{\tilde{n}_{x}}(m^{2})}\\ =\sum\limits_{\begin{subarray}{c}m=(m^{1},m^{2})\in\mathbb{N}_{0}^{E\cup E_{g}}:\\ m^{1}_{\{x,g\}}=m^{2}_{\{x,g\}}=\tilde{n}_{x}(m^{1})=\tilde{n}_{x}(m^{2})=0\,\,\forall x\in A\\ m_{\{x,g\}}^{1},m_{\{x,g\}}^{2}\in r\mathbb{N}_{0}\,\,\forall x\in V\end{subarray}}\prod_{i=1}^{2}\prod_{e\in E}\frac{\beta^{m^{i}_{e}}}{m^{i}_{e}!}\prod_{e\in E_{g}}\frac{(\beta h)^{m^{i}_{e}/r}}{(m^{i}_{e}/r)!}\\ \times\int_{\Omega_{s}}\boldsymbol{ds}\,\prod_{z\in V}\,{\gamma_{z}(\boldsymbol{s})}\,\prod_{x\in V}{(S_{x}^{1})}^{{\tilde{n}_{x}}(m^{1})+m^{1}_{\{x,g\}}-m^{2}_{\{x,g\}}}{(S_{x}^{2})}^{{\tilde{n}_{x}}(m^{2})}\\ =\sum\limits_{\begin{subarray}{c}m\in\Sigma_{r}\,:\\ m^{1}_{\{x,g\}}=m^{2}_{\{x,g\}}={\tilde{n}_{x}}(m^{1})={\tilde{n}_{x}}(m^{2})=0\,\,\forall x\in A\\ m^{1}_{\{x,s\}}=m^{2}_{\{x,s\}}=0\,\,\forall x\in V\end{subarray}}\prod_{i=1}^{2}\prod_{e\in E}\frac{\beta^{m^{i}_{e}}}{m^{i}_{e}!}\prod_{e\in E_{g}}\frac{(\beta h)^{m^{i}_{e}/r}}{(m^{i}_{e}/r)!}\prod_{x\in V}U({\tilde{n}_{x}}(m^{1})+m^{1}_{\{x,g\}}-m^{2}_{\{x,g\}}),$ (8.11)
where for the first identity we applied (8.4) at each vertex $x\in A$ to observe that the integral equals zero if there exists some edge $\{x,y\}$ with $x\in A$ and $y\sim x$ or $y=g$ with $m_{\{x,y\}}^{i}>0$ for some $i\in\{1,2\}$, for the second step we used (8.5) at each vertex $x\in V\setminus A$, in analogy to (8.9), obtaining the analogous of (8.10). We then obtained the numerator in the RHS of the first identity in Proposition 3.12, thus concluding the proof.
We now move to the proof of the second claim in the proposition by setting for $p\in\mathbb{N}_{0}$, $f=\prod_{z\in A}\frac{\overline{(S_{z}^{1}S_{z}^{2})}^{p}}{(2\pi)^{2}\gamma_{z}(\boldsymbol{s})}.$ We obtain from (8.7) that, under this choice of $f$, (8.6) equals
$\sum\limits_{\begin{subarray}{c}m=(m^{1},m^{2})\in\mathbb{N}_{0}^{E\cup E_{g}}:\\ m_{\{x,g\}}^{1},m_{\{x,g\}}^{2}\in r\mathbb{N}_{0}\forall x\in V\end{subarray}}\prod_{i=1}^{2}\prod_{e\in E}\frac{\beta^{m^{i}_{e}}}{m^{i}_{e}!}\prod_{e\in E_{g}}\frac{(\beta h)^{m^{i}_{e}/r}}{(m^{i}_{e}/r)!}\\ \times\int_{\Omega_{s}}\boldsymbol{ds}\,\Big(\prod_{z\in V\setminus A}\,{\gamma_{z}(\boldsymbol{s})}\Big){\Big(\frac{U(p)}{{(2\pi)}^{2}}\Big)}^{|A|}\prod_{x\in V}{(S_{x}^{1})}^{{\tilde{n}_{x}}(m^{1})+m^{1}_{\{x,g\}}-m^{2}_{\{x,g\}}-p\mathbbm{1}_{A}(x)}{(S_{x}^{2})}^{{\tilde{n}_{x}}(m^{2})-p\mathbbm{1}_{A}(x)}$ (8.12)
where $\mathbbm{1}_{A}(x)=1$ if $x\in A$ and $\mathbbm{1}_{A}(x)=0$ otherwise. By using (8.5) we deduce that, for each $m=(m^{1},m^{2})\in\mathbb{N}_{0}^{E\cup E_{g}}$ in the previous sum,
$\int_{\Omega_{s}}\boldsymbol{ds}\,\Big(\prod_{z\in V\setminus A}\,{\gamma_{z}(\boldsymbol{s})}\Big){\Big(\frac{U(p)}{(2\pi)^{2}}\Big)}^{|A|}\prod_{x\in V}{(S_{x}^{1})}^{{\tilde{n}_{x}}(m^{1})+m^{1}_{\{x,g\}}-m^{2}_{\{x,g\}}-p\mathbbm{1}_{A}(x)}{(S_{x}^{2})}^{{\tilde{n}_{x}}(m^{2})-p\mathbbm{1}_{A}(x)}\\ =\prod_{x\in V\setminus A}U({\tilde{n}_{x}}(m^{1})+m^{1}_{\{x,g\}}-m^{2}_{\{x,g\}})\,\delta_{{\tilde{n}_{x}}(m^{1})+m^{1}_{\{x,g\}},{\tilde{n}_{x}}(m^{2})+m^{2}_{\{x,g\}}}\\ \times\prod_{x\in A}U(p)\,\,\delta_{{\tilde{n}_{x}}(m^{1})+m^{1}_{\{x,g\}},p}\,\,\delta_{{\tilde{n}_{x}}(m^{2})+m^{2}_{\{x,g\}},p}.$ (8.13)
The proof now follows the same lines as the proof of Proposition 3.11 after replacing the previous expression in (8.12). ∎